### v1

# Force Fields for Deep Eutectic Mixtures: Application to Structure and 2D-Infrared Spectroscopy

Abstract
Parametrizing energy functions for ionic systems can be challenging. Here, the total energy function for an eutectic system consisting of water, SCN-, K+ and acetamide is improved vis-a-vis experimentally measured properties. Given the importance of electrostatic interactions, two different types of models are considered: the first (model M0) uses atom-centered multipole whereas the other two (models M1 and M2) are based on fluctuating minimal distributed charges (fMDCM) that respond to geometrical changes of SCN-. The Lennard-Jones parameters of the anion are adjusted to best reproduce experimentally known hydration free energies and densities which are matched to within a few percent for the final models irrespective of the electrostatic model. Molecular dynamics simulations of the eutectic mixtures with varying water content (between 0 and 100 %) yield radial distribution functions and frequency correlation functions for the CN-stretch vibration. Comparison with experiments indicate that models based on fMDCM are considerably more consistent that those using multipoles. Computed viscosities from models M1 and M2 are within 30 % of measured values and their change with increasing water content is consistent with experiments. This is not the case for model M0.
si University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
1 Introduction
Characterizing the energetics, structural dynamics and spectroscopy of
heterogeneous systems in the condensed phase is challenging from a
computational and an experimental perspective. Typical examples of
such endeavours are hydrated proteins in a representative environment
including the necessary anions and cations, ionic liquids, or eutectic
mixtures. From a computational perspective the main challenges are the
accuracy of the energy function that is required to carry out
atomistic simulations and the time scale on which such simulations can
be run. On the other hand, from an experimental viewpoint, one of the
greatest challenges is the fact that usually the “full system” is
probed which may lead to spectral congestion if optical spectroscopy
is used as an example. Also, covering multiple time scales from the
femtosecond to the second time scale is demanding, although recent
progress has been made here.1
For meaningful atomistic simulations sufficiently accurate energy
functions are a prerequisite. Although there has been impressive
progress in directly solving the electronic Schrödinger equation for
increasingly large systems, doing so along a molecular dynamics
simulation is usually unfeasible beyond the multiple-ten-picosecond
time scale, in particular for methods superior to density functional
theory. On the other hand, empirical energy functions (“force
fields”) are computationally advantageous to evaluate. Propagating
the dynamics based on Newton’s equations of motion allows to reach
long time scales for large molecular systems, e.g. sub-microsecond
for atoms.2 Such molecular dynamics (MD)
simulations have been used to investigate processes ranging from
protein folding,3 ligand binding,4
crowding in cellular environments5 to characterizing
spectroscopic properties of solutes and peptides and reactions in the
gas phase and in solution.6 However one of the
challenges remains to develop energy functions that retain the
precision of the QM methods they are often based on.
One- or multi-dimensional vibrational spectroscopy is a powerful means
to characterize the structural dynamics of complex
systems.7, 8 The molecular-level understanding
and interpretation of the spectral signatures for such systems is
greatly aided by a combined simulation/experiment
approach.9, 10 The atomistic simulations
often rely on energy functions that parametrize the bonded and
nonbonded interactions. Traditionally, empirical energy functions use
harmonic springs for chemical bonds and valence angles, periodic
functions for dihedrals, an atom-centered point charge-based model for
charges and a Lennard-Jones representation for van der Waals
interactions together with additional, more purpose-tailored
terms.11 For applications to spectroscopy the
chemical bonds (stretching vibration) need to be described somewhat
more realistically to account for mechanical anharmonicity for which
Morse-oscillators are often sufficient because experiments at ambient
temperatures are not sensitive to highly vibrationally excited
states. Nevertheless, there is scope to use more accurate
representations, for example based on machine learning-type
approaches, in particular if reference data from high-level electronic
structure calculations are available. For the nonbonded interactions -
which include electrostatic and van der Waals terms - more
physics-based models that go beyond the standard representations have
been developed.
The first-order treatment of the electrostatic interaction is based on
atom-centered point charges for Coulomb interactions. Such pair
interactions can be rapidly computed but lack the accuracy for
describing anisotropic contributions to the charge
density.12 Including higher-order atomic multipoles
improves the accuracy but at the expense of increased computational
cost and implementation
complexity.13, 14, 15, 16
Accounting for polarizability is another contribution that has been
recently included in empirical force fields and shows much promise for
further improvements of the computational models.17 From
an empirical force field perspective the van der Waals interactions
are often represented as Lennard-Jones terms with ad hoc
(Lorentz-Berthelot) combination rules. Alternative and potentially
improved representations are the buffered 14-7
parametrization18 and modified combination
rules19, 20
Deep eutectic mixtures (DES) are electrostatically demanding systems.
The present work aims at parametrizing atomistic force fields using
state-of-the art methods by combining machine learning-based
approaches for bonded and nonbonded terms, refinement of the
Lenard-Jones interactions with respect to thermodynamic data and
validation on structural, spectroscopic and thermodynamic
measurements. First, the methods are presented, followed by the
reparametrization and validation of the energy functions. Next,
extended MD simulations are analyzed with respect to pair distribution
functions and the frequency fluctuation correlation functions from
experiment and simulations are compared. This is followed by a
discussion and conclusions.
2 Computational and Experimental Methods
2.1 Simulation Setup
Molecular dynamics simulations were performed using the CHARMM
program.21 The molar composition of the systems
was changed by modifying the number of water and acetamide molecules
while keeping constant the total concentration of K+ and
SCN-. The number of molecules and molar fractions are reported in
Table S1. Bonds involving hydrogen atoms were
constrained using the SHAKE algorithm.22 For the nonbonded
interactions the cutoff was 14 Å and electrostatic interactions
were treated using the Particle Mesh Ewald algorithm.23
In total, 5 independent random initial configurations for each of the
9 system compositions were set up using PACKMOL.24
Following 100 ps of heating and equilibrium simulation, each,
production simulations in the NpT ensemble at 300 K and normal
pressure (1 atm) were performed for 5 ns with a time step of 1 fs
using the leap-frog integrator and a Hoover
thermostat.25 For each system composition a total of
ns were sampled.
2.2 Inter- and Intramolecular Interactions
The heterogeneous system considered in the present work (water,
acetamide, K+, SCN-) was described by a combination of
parametrizations based on the all-atom force field CHARMM36
(CGenFF)26, the corresponding TIP3P water
model27 to be used together with CGenFF, and
optimized parameters for the ions. Two sets of nonbonding
van-der-Waals (vdW) parameters for the potassium cation K+ with an
assigned atom charge of were adopted from previous
work28, 29 and acetamide was described by
CGenFF. Dedicated parametrizations for bonding and nonbonding terms
were used for the thiocyanate anion (SCN-) as described next.
Bonded Interactions: The bonding potential for the thiocyanate
anion (SCN-) was a reproducing kernel Hilbert space
(RKHS)30, 31 representation based on ab
initio data at the PNO-LCCSD(T)-F12/aug-cc-pVTZ-F12 level of theory
using the MOLPRO program package.32, 33 The
geometry of SCN- was described in terms of Jacobi coordinates
where is the CN separation, is the distance
between S and the center of mass of CN and is the angle
between the vectors and . Reference energies were
determined on an equidistant grid defined by Å (8 points), Å (9 points),
and (6 points). This yields a total
of 432 grid points. Energies for grid points within kcal/mol of the equilibrium structure
( Å, Å and
) were retained for constructing the
RKHS. The root mean squared error (RMSE) between reference and the
RKHS representation was kcal/mol. Two-dimensional cuts
through the RKHS representation for SCN- along Jacobi distances
and for different angles are shown in Figure S1.
Electrostatic Interactions: For the electrostatic interactions
the electrostatic potential (ESP) of SCN- was represented as a
flexible minimally distributed charge model
(fMDCM).34 For this, 8 point charges were
distributed around the SCN- atoms with positions within the local
axis frame determined by 3rd order polynomial functions with
and the SCN- bond angle . The 96
parameters of the 24 polynomial functions - 4 parameters per
polynomial for each Cartesian coordinate (3) and distributed charge
(8) - were optimized to best reproduce the reference ESPs for
different SCN- conformations. The reference ESPs on a range of
geometries were determined at the M06-2X/aug-cc-pVTZ level of theory
using the Gaussian suite of programs.35 The grid
contained 289 geometries, including 32 SCN valence angles (]) and 3 bond lengths for the SC ( Å) and CN ( Å) separations.
Starting from the initial 8 charge MDCM model the reference ESP was
fit for SCN- in a slightly bent conformation ()
according to the optimization scheme presented in
Refs. 36, 37. Starting from this model as
an initial guess for the polynomial parameters, the quasi-Newton BFGS
minimization method38 was used to optimize
the polynomial parameters by minimizing the RMSE between reference and
computed ESP grid point values for each conformation. The total RMSE
was the sum of the RMSE per conformation weighted by a factor to prioritize an accurate ESP fit for
low-energy conformations. Here, was the ab initio energy
of SCN- in conformation relative to the energy of the
equilibrium conformation . Furthermore, the 8
distributed charges were constrained to lie only along the bond axis
for linear SCN- conformations, the charge displacement from the
closest reference atom and its charge magnitude within a0 and respectively. Figure
1A shows the the distributed charge position predicted
by the fMDCM model for SCN- in linear equilibrium conformation
where all charges a aligned along the SCN- axis. For angled
SCN- conformation such as in Figure 1B with a bond
angle of 160∘ the charge positions deviates from the bond
axes. Note the two negative distributed charges (red spheres) closest
to the sulfur atom (transparent yellow sphere) which are merged into
the same position for linear SCN-, see Figure 1A, and
become symmetrically lifted and mirrored at the SCN- plane in
Figure 1B.
The RMSE between the fMDCM model and the reference ab initio
ESP was kcal/mol on average and kcal/mol for the
linear equilibrium structure of SCN-. Following the
literature,39 the diagonal elements of computed
polarizability for SCN- in the gas phase
(M06-2X/aug-cc-pVTZ) were Å3 and
Å3 to yield an average isotropic
polarizability of Å3 However,
computational studies have shown a reduction in the polarizability for
anions from gas phase to water solution by between 5% and 34%,
respectively.40, 41 Thus, a reduced
polarizability of Å3 was equally distributed to the atoms
of SCN- with Å3 each.
In addition to the new fMDCM model, a multipolar
(MTP)14, 42, 43 representation of
SCN- used in previous work was used as well.10
The electrostatic atomic moments up to the quadrupole moment were
optimized to fit the ESP of SCN- in its equilibrium conformation
computed at the MP2/aug-cc-pVDZ level of theory.
Van-der-Waals Interactions: Changing the electrostatic model for
SCN- from point charges to a more elaborate fMDCM or MTP model
also requires to readjust the Lennard-Jones
parameters.44 Three models were considered in the
following for SCN-. Model (M0) was based on MTP with LJ
parameters originally from Bian et al.28 and
scaled to best reproduce measured thermodynamic properties. Model M1 used fMDCM electrostatic instead of MTP and initial LJ
parameters from Tesei et al.,29 and model M2
was a combination of fMDCM and optimized LJ parameters with respect to
reference ab initio interaction energies as described
below. Experimentally observed density of aqueous KSCN solution and
the hydration free energy of SCN- in water were then determined
for each of the models M0, M1 and M2. For the K+
cation the LJ parameters were those from the respective parameter set
in the literature (Ref. 28 for M0 and M2
and Ref. 29 for M1) and for model M2 the
LJ parameters from Bian and coworker were chosen for
K+.28
2.3 Analysis
Hydration free energies for the SCN- anion in water solvent were computed from thermodynamic integration.9 One SCN- anion was sampled in the gas phase and in pure water. The condensed-phase simulations were carried out in the NpT ensemble with 997 water molecules (cubic box size Å3).45, 9 Simulations with 24 different coupling parameters for electrostatic and vdW interactions, respectively, were performed for each a gas phase and solvated SCN-. Initial conditions for these simulation were taken from an unbiased simulation, equilibrated for 50 ps with the respective coupling parameter and run for another 150 ps for statistical sampling. In general, the hydration free energy is accumulated from
For a triatomic such as SCN-,
due to the 1-2 and 1-3
nonbonded interaction exclusion.46 Therefore, only
and
needed to be accumulated.
The density of an aqueous KSCN solution was determined from 500 ps
simulations in the ensemble. The system composition corresponds
to the 100% one reported in Table S1 with a KSCN
molality of mol/kg. The equilibrium
simulation box volume is computed as the average box volume within the
last 100 ps of the simulations (400–500 ps in the production run).
Convergence within the reported precision was checked by comparing
with the average taken from the results of the full production run of
500 ps.
For computing the frequency fluctuation
correlation function (FFCF), the frequency trajectories
for each oscillator are required. These were determined from an
instantaneous normal mode (INM)47 analysis for the CN
vibrational frequencies for all SCN- ions
on snapshots of the trajectories every fs along
the first 2 ns of each production simulation (10 ns in total for
each system composition).
For the INM analysis the structure of every SCN- ion was optimized
while freezing the position of all remaining atoms in the system,
followed by a normal mode analysis using the same force field that was
employed for the MD simulations. Previously, such an approach has been
validated for N in solution by comparing with rigorous quantum
bound state calculations.48
From the frequency trajectory for each oscillator the FFCF was determined which contains information on relaxation time scales corresponding to the solvent dynamics around the solute. The FFCFs were fit to an empirical expression
using an automated curve fitting tool from the SciPy
library.49 Here, , , and
are the amplitudes, decay time scales, phase and
asymptotic value of the FFCF.
2.4 2D-IR Spectroscopy
The experimental data shown in this paper (Figure 7) originate from the same raw data as those of Ref. 10, this time extracting the FFCF from the 2D-IR spectra. In brief, a 2D IR spectrometer in pump-probe with scatter suppression by quasi-phase-cycling has been used to obtain the raw data.50 A 50%/50% mixture of two isotopologues of the salt, KS12C14N and KS13C15 has been investigated. Solutions were prepared with constant total salt concentration of 4.4 M, and the molar ratio of D2O vs acetamide varied from 0% to 100% in steps of 10%.
To extract the essential information from the 2D IR spectra, i.e., the FFCF, each 2D IR peak was modelled as a correlated 2D-Gaussian function:
where is an amplitude, the center frequencies along
frequency axis , the frequency width, and a
correlation coefficient.51 In total 8 such peaks were
defined for a 2D IR spectrum, with many parameters set equal to
minimize their total number in the global fitting. That is, two pairs
of diagonal peaks were defined for the two isotopologues with
different central frequencies and amplitudes , but
identical width and correlation , since the latter
two parameters describe the identical interaction of the ions with the
solvent environment. Each diagonal peak consists of ground state
bleach (GB)/stimulated emission (SE) as well as excited state
absorption (EA), where the latter has been down-shifted in the
-direction by the anharmonicity
. The amplitudes and correlations of
the two contributions to each diagonal peak were assumed to be the
same. In addition, two cross peak pairs were introduced, each again
consisting of GB/SE and EA contributions, for which center frequencies
and width were assumed to be the same as for the corresponding
diagonal peaks along the two frequency axes. The amplitudes of the
two cross peak pairs were set to be the same, and was assumed
for the correlation, since cross-relaxation randomizes the
frequency.
The property of interest here is the frequency correlation of the diagonal peaks as a function of the waiting time between the pump-pulse pair and the probe pulse. For times larger than a few 100 fs, reflects the FFCF and can be compared to results from MD simulations after proper scaling, while dynamics faster than this is hidden in the homogeneous dephasing time .8
3 Results
3.1 Adaptation and Validation of the LJ Parameters
Three models M0 to M2 with different treatment of the
electrostatics (MTP or fMDCM) and corresponding LJ parameters for
SCN- and K+ were considered. Details on the models are given in
Tables S2 to S4. The
experimental reference data required for models M0 and M1
were the density of aqueous KSCN solution
( g/cm3) at the equivalent KSCN molality
of mol/kg52 and the
hydration free energy . For SCN- an
estimated value of kcal/mol is
available53 which compares with related anions such as
HS- ( kcal/mol), N ( kcal/mol)
54, or CN- ( kcal/mol)55.
For model M0 using the MTP electrostatic model, the
literature28 LJ parameters of the
SCN- atoms were scaled by multiplying with a factor . Using these scaled parameters, the
system density and hydration free energy of SCN- were determined,
respectively. Figure 2 (blue lines) reports the
variation of and with scaling and
provides an estimate for to best reproduce the two
experiments. With the average density is
g/cm3 and the hydration free energy of
SCN- in water is kcal/mol. These are significant improvements of 10% and 20%,
respectively, over the two reference values for (i.e. using
the original LJ parameters).
Model M1 (fMDCM electrostatics) was based on LJ parameters that
had been used for modeling aqueous SCN- solutions.29
To adjust the literature LJ parameters to the fMDCM electrostatic
representation, values - which are equivalent to
by way of - of the
SCN- atoms were again scaled by a factor to best
reproduce the experimentally measured observables. Figure
2 (green lines) reports and for different scaling factors in a scan range of
and the experimental reference (horizontal black dotted
and dashed lines). For model M1 and , the
computed density g/cm3 compares well
with the experimentally measured52 density
g/cm3 and the computed hydration free
energies kcal/mol
( kcal/mol for unscaled LJ parameter) is within reasonable
range of the experiments.
For model M2, also based on the fMDCM representation for the
SCN- anion, a different approach was taken. Here, the LJ parameters
and were adjusted to best match
counterpoise-corrected interaction energies from electronic structure
calculations. The systems considered included on SCN- anion
surrounded by (i) 16 water molecules; (ii) 14 water molecules with
either 1 SCN- or (iii) 1 K+ ion; (iv) 12 water molecules with
both 1 SCN- and K+ ion. For each type of system (i) to (iv) 50
independent conformations were extracted randomly from previous MD
simulations.10 Reference interaction energies were
then determined at the M06-2X/aug-cc-pVTZ level of theory using
Gaussian program.35 Higher levels of quantum chemical
theory, such as coupled cluster energies, are not feasible due to the
unfavourable scaling of the computations with system and basis set
size.
The interaction energy consists of electrostatic contributions using
the fMDCM model for SCN- and the vdw interactions described as a LJ
potential. The optimized LJ parameters achieved an overall RMSE of
kcal/mol for all snapshots. The individual RMSE values are (i)
kcal/mol, (ii) kcal/mol, (iii) kcal/mol, and
(iv) kcal/mol. The reference interaction energies cover a broad
range from kcal/mol to kcal/mol, see Figure
3. In a second step, the optimized of
the SCN- atoms were again scaled by a factor and the density of
an aqueous KSCN solution and hydration free energy were computed from explicit MD simulations, see
Figure 2 (red lines). Closest agreement with
experiment is found for which yields
g/cm3 and kcal/mol. Hence, fitting the
interaction energies for a diverse set of hydrated structures yields
LJ parameters that are suitable for thermodynamic observables from
simulations.
In conclusion, all three models allow to adjust the Lennard-Jones
parameters for MTP and fMDCM electrostatics such as to satisfactorily
reproduce experimentally measured thermodynamic reference data, see
Figure 2. The actual scaling parameters are close to
for the fMDCM models (, ) and larger by % for the van der Waals ranges if MTP
() is used. Scaling factors of for
MTP-based models are consistent with earlier work on the vibrational
relaxation and hydration free energies of CN- and
OH-.9
3.2 Structure and Ordering of the Mixture
Using the optimized LJ parameters for models M0 to M2,
equilibrium MD simulations 25 ns in length were carried out for each
system composition. From these simulations, radial pair distribution
functions were determined and compared with results from
empirical potential structure refinement (EPSR) fits to best reproduce
neutron diffraction measurements of aqueous KSCN
solution.56 It is important to stress that such pair
distribution functions are not measured directly from experiments but
rather adjusted empirically to reproduce the neutron scattering
amplitude which is written as a weighted average over pair
distribution functions.56 Figure 4
reports SCN-–water (top) and SCN-–SCN- (bottom) pair
correlation functions for different mixtures KSCN in solution (0,
50%, 80%, 100% water content) compared with experimentally obtained
. The radial distribution functions in Figures
4B and C (models M1, M2) are considerably
smoother than those using the MTP model in Figure 4A
(M0). For simulations in pure water (100%) direct comparisons
between simulations (dot-dashed black line) and EPSR results (solid
black line) are possible.56 For model M0, the
overall shape of agrees reasonably well with the
EPSR results although the computed pair distribution function is
overstructured with a local maximum around 5.5 Å. Such maxima are
not found for models M1 and M2 for which the position of
the maximum is shifted to larger separations and the height of the
maximum is overestimated for M1 but quite well described by M2. With decreasing water content the height of the first peak
increases - i.e. the ion recruits water molecules as the water density
decreases. The shift of the first maximum for model M2 relative
to model M1 is consistent with the larger values
for the S- and N-atoms for M2 compared with M1, see Tables
S3 and S4.
For the C–OW and S–OW pair correlation functions
and , similar observations are
made. In all cases, the pair distribution functions from model M0 are overstructured with additional peaks at larger separations
(particularly for ). On the other hand, radial
distribution functions from simulations using models M1 and
M2 are in reasonably good agreement with EPSR results - in
particular for large separations - whereas for the first solvation
shell the position of the first maximum is at larger separations and
the peak heights for M2 are closer to the EPSR results.
The difference in the peak intensities for obtained from the
MTP and fMDCM simulation indicates favouring considerably stronger
binding of H2O to SCN- ions by the MTP model. This also leads to
a more rigid second solvation shell indicated by significant peak
heights Åat larger values of and is
consistent with the increased hydration free energies even after
scaling the values by , see Figure
2.
For simulations of KSCN in water/acetamide solvent mixtures with fMDCM
(models M1, M2), the peak intensities in between
SCN- atoms and water oxygen increases for larger acetamide
concentration (with exception of the water-free system 0%) but the
peak positions remain unchanged. A similar observation for different
mixtures is made for simulation with model (M0), but with
considerably more pronounced increases in the peak intensities with
larger acetamide content in the solvent mixture.
For anion-anion pair distribution functions (Figures 4D
to F) the agreement for for models M1 and M2 is particularly notable. Notwithstanding, model M0 is also
capable of qualitatively describing the experiments. For the S–N and
N–N pair correlation functions model M0 has the first maximum
either shifted to longer separations or a much too pronounced
maximum. Compared with that, models M1 and M2 correctly
capture the peak height, which is a proxy for the interaction
strength, whereas the position of the first maximum is in general
shifted to shorter separations. This, however, changes with decreasing
water content for which the S–N and N–N pair correlation functions
shift to longer separations. It should be noted that the peak position
of the anion-anion pair distribution functions in simulations can be
steered by adjusting the LJ parameter of the SCN-
atoms individually. However, the procedure followed here was not to
match pair distribution functions as the LJ parameters of
SCN- are at first either taken from literature (M0, M1)
or fit to ab initio data (M2) and later scaled
collectively to best match the density of an aqueous KSCN system or
hydration free energy of SCN- in water.
The comparison of the between atoms of different SCN- anions
in simulations using the three models for the intermolecular
interactions are shown in Figures 4D to F. Peak
intensities for SCN- atom pair in different water/acetamide solvent
mixture remains either similar (M1) or increase for larger
acetamide content (M2). For simulations with multipoles (model
M0), the peak positions are generally closer those reported from
experiments but with higher intensities which are highest for pure
water solvent and decreases for larger acetamide content in the
water/acetamide solvent mixture.
Another view on the structural SCN- radial-angular pair
distribution is shown in Figure 5 along the C-C bond
distances and (CN)-(CN) bond angles (similar to SCN- dipole
vectors). Even after adjusting the LJ parameters for M0 to yield
improved thermodynamics, the MTP-based model lead to mainly angled
SCN- pairs in pure acetamide and solvent mixtures with large
acetamide content. This changes to predominantly parallel and, partly,
anti-parallel SCN- pair alignments with increasing water content
all of which is consistent with results from simulations using the
unmodified MTP-based model.10 In simulations using
fMDCM for the electrostatics (models M1 and M2), the
angular pair SCN- distribution functions depend less strongly on
solvent mixtures and are close to squared cosine angle distributions
which are statistically expected for weakly or non-interacting
particles. This can be even better seen in Figure 6
which reports the 2-dimensional projections of the distributions shown
in Figure 5 using model M2.
Contrary to M0, the probability distributions do not change so
profoundly with increasing water content for models M1 and M2. Nevertheless, the centroid of the C–C probability
distributions shift to somewhat longer C–C separations with
increasing water content. Also, the maximum for
is more sharply peaked for low water content and broadens around the
maximum for pure water as the environment. Conversely, the two
higher-probability regions at large C–C separations Å
that appear for acetamide as the solvent disappear as the water
content increases.
3.3 Dynamics of the Mixtures
Finally, the dynamics of the mixtures was studied by analyzing the
C–N vibrations from instantaneous normal modes of the SCN- anion
and the corresponding frequency fluctuation correlation functions
(FFCF) for varying water content. Figure 7A to C shows
the FFCFs with models M0 to M2, respectively. Figure
7D reports the experimentally measured FFCFs from the
2D-IR experiments. It is found that the water content affects the
FFCFs in a pronounced fashion: with increasing water content the fast
and slow time scales decrease. For the FFCFs from the simulations a
bi-exponential fit is a meaningful model applicable to all
mixtures. Assuming a single exponential significantly deteriorates the
fit. To allow direct comparison between experiment and simulations,
the experimental data was fit to a double-exponential decay because
for the computed FFCFs a second, short-time decay (to which the
experiments are insensitive) needs to be included for fitting the full
time range of the data, see Figures S2A to C. The
set of optimized parameters of the bi-exponential fit function are
shown in Figure S3 and Table S6.
Qualitatively, models M1 and M2 are rather successful in
capturing the experimentally observed FFCFs and their dependence on
the amount of water in the solvent mixture whereas model M0 is
not. Overall, the decay from simulations using M0 is too
slow. However, the relative amount by which the time constant changes
between 0% and 100 % water is considerably closer to the
experimentally observed change than in the other two models.
A more quantitative analysis is based on the decay times of the FFCFs
depending on water content, see Figure
8. Experimentally, the slow decay (black filled triangles) almost monotonously decreases with
increasing water content. This is qualitatively reproduced from
simulations using all three models although the amplitude from model
M0 (green filled diamonds) is larger by a factor of and
those from models M1 and M2 (red filled circles and blue
filled squares) are larger by a factor of and the slope is
considerably steeper. For the fast decay
(black open triangles) an increase with increasing water content up to
% is observed followed by a monotonous decrease. This is
qualitatively captured by model M0 (green open diamonds),
although the maximum is shifted to higher water content, whereas
models M1 and M2 (red open circles and blue open squares)
again find monotonous decrease to amplitudes considerably lower than
the experiment for pure water.
However, it is known that in biexponential fits the final parameters
can be strongly correlated. Therefore, it may be more appropriate to
compare, for example, amplitude-weighted decay times , see
Figure 8B, with amplitudes reported in Figure
S3A. Overall, the computations reproduce the
decrease in with increasing water content. For M0 the experimentally determined values are typically overestimated
by an order of magnitude or less. This differs for models M1 and
M2 which overestimate for low water content
and underestimate it for high water content. Specifically, for the
water-free system (0 %) ps which
compares with ps from
models M1 and M2 and ps. Together with the results from Figure 7 it is
concluded that the two fMDCM-based models more realistically describe
the solvent dynamics as encapsulated in the FFCFs.
4 Discussion and Conclusion
The present work improves, validates and applies an energy function
for a heterogeneous eutectic mixture consisting of water, ions and
acetamide. For the SCN- ion a flexible MDCM model was
developed. The van-der-Waals parameters were optimized vis-a-vis
experimental densities and hydration free energies using the TIP3P
water model for consistency with the remaining force field
parameters. This point will be considered further below. The bonded
terms for SCN- were replaced by an accurate reproducing kernel
representation of reference calculations at the
PNO-LCCSD(T)-F12/aug-cc-pVTZ-F12 level of theory. Three different
parametrization approaches were pursued and all of them describe the
experimental reference data satisfactorily.
Force fields for deep eutectic solvents have been previously optimized
using different protocols and
approaches57, 58, 59, 60, 61, 62, 63, 64
Most common to these efforts is to start from a conventional energy
function such as the General Amber Force Field (GAFF),
65 adapted for particular
applications66 and to further readjust
parameters.57, 63 The adjustment is usually
to scale partial charges62 and in some cases
van-der-Waals parameters to reproduce observed properties such as
diffusivities, viscosities or the densities.63 In terms
of using diffusivities for the parametrization from explicit MD
simulations it is noted that in DES they are several magnitudes slower
than those in typical liquids such as water ( to
m2/2 vs.
m2/s).61, 57, 67 Thus the
determination of the values typically require a set of ns simulations (compared with 5 ns for water67)
which would make an explicit parametrization extremely expensive.
Alternatively, there have been efforts to develop models based on
symmetry adapted perturbation theory (SAPT)61 which
were refined by comparing with first principles MD
simulations. Another approach was chosen for the Canongia Lopes &
Padua (CL&P) force field and its polarizable variant
CL&Pol.58 Recent fine-tuning efforts used structural
data.59 The present status of computational approaches
for deep eutectic solvents has been recently
reviewed.64
In the present work two alternative procedures were followed, both
based on improved electrostatics for the SCN- anion and
subsequently adjusting the LJ parameters. First, a previously used
distributed multipole model was improved by scaling the LJ parameters
to correctly describe experimentally measured liquid densities and
hydration free energies (model M0). To probe the influence of
the electrostatic model used, a fluctuating MDCM model was constructed
to replace the multipoles, and the LJ parameters were adjusted as for
the multipolar model (model M1). Second, reference interaction
energies from hydrated SCN- clusters of different compositions
were determined from electronic structure calculations (model M2). After subtracting the electrostatic contributions (fMDCM for
SCN-) the LJ parameters were scaled to best describe the
remaining nonbonded interactions. Finally, a second round of
refinement by scaling the LJ parameters of SCN-
followed to best describe again experimental liquid densities and
hydration free energies as for M0 and M1.
Interestingly, the optimal scaling parameter for model M2 to
best describe experimental observables was . This suggests that fitting LJ parameters to interaction energies
from sufficiently high-level electronic structure calculations and
using a robust charge representation for the ion (fMDCM) yields models
that can be used for meaningful simulation of thermodynamic
observables. It will be interesting to see how generic such an
approach is by applying it to other, similar systems. On the other
hand, model M0 required rather large changes in the LJ
parameters (up to 10 %) to obtain agreement with the thermodynamic
data considered here. This is consistent with earlier work which also
found that in particular LJ parameters (or
)) needed to be increased by % for hydration free energies consistent with
experiment.9
Decomposition of a molecular ESP into atomic distributed multipoles
can be viewed as a linear combination of atom-centered spherical
harmonics. The expansion coefficients are the atomic charge, dipole,
and higher order multipole moments. Usually, this series is truncated
at the atomic quadrupole. The basis functions of this expansion
(“atomic orbitals”) have a rather pronounced directionality which
influences short- to mid-range intermolecular interactions. Contrary
to that, point charge-based models such as fMDCM represent the
molecular ESP as a superposition of isotropic potentials. This can be
expected to lead to less pronounced angular variations. Both models -
multipole- and PC-based - can be fit to the same quality compared with
reference molecular ESPs from electronic structure
calculations. Qualitatively, the increased LJ
parameters required for model M0 (MTP-based) compared with M1 (PC-based) indicate that the more pronounced anisotropies from
atom-based multipoles need to be damped to moderate interactions with
nonbonded partners and to potentially reduce the pronounced
anisotropy.
The changes in the LJ parameters for model M0 also influence
measured spectroscopic properties. For example, the vibrational energy
transfer between neighboring SCN-, which was measured
previously,10 is affected by intermolecular
coupling and structural alignment. The more pronounced anisotropy of a
MTP-based representation is also seen in the 3-dimensional pair
distribution functions , see Figure
5. For model M0 the angular dependence of these
distribution functions is considerably more structured compared with
models M1 and M2. It is also interesting to note that
despite increasing the LJ by % the
for M0 from the present work
compared with the earlier employed
parametrization10 do not differ appreciably
(compare with Figure 5 in Ref.10)
Using the three interaction models the viscosities for varying
acetamide/water ratios were determined as an additional validation,
see Figure S4. For this, the stress tensor
was computed and the viscosity was determined from
where
are the upper triangular elements of the modified stress tensor
.68, 63 Due to the
strong intermolecular interactions and high viscosities, converging
can be rather demanding and is not attempted here. Rather, 5
independent simulations were carried out for each composition
and the results were averaged to obtain illustrative results.
For model M0 the viscosity first fluctuates around
mPas before decaying with increasing water content (Figure
S4A). On the other hand, models M1 and M2 find that monotonously decreases with increasing water
content, see Figures S4B and C. The maximum
viscosities, which occur for zero water content, differ by a factor of
two. Experiments for the water-free system report
mPas at 303 K.69 Simulations with M1 and
M2 are consistent with such values. More importantly, however,
the experiments report that decreases monotonously with
increasing water content what is found by models M1 and M2
but not M0. The magnitude of for the three models also
reflects the strength of the intermolecular interactions. Given the
large scatter in from individual simulations, in particular for
low water-content, reaching convergence requires longer or more
simulation data. From a computer efficiency perspective it appears to
be advisable to consider as a validation property rather than a
property used for fitting the intermolecular interactions.
As a final comment, the TIP3P water model used here is known to have
various limitations when comparing experimentally measured quantities
with those from simulations.70 As was already
mentioned, the reason to use the TIP3P model was consistency with the
CGenFF parameters for the other species present in the eutectic
mixture. To quantitatively assess changes using a different water
model, MD simulations were carried out using model M2 but with
the TIP4P water model.27 Again, hydrated
SCN- clusters were extracted, reference interaction energies
computed, and the parametrization step was repeated as had been done
for TIP3P, see Figure 3. The results are reported in
Figure 9. It is found that the two water models yield
similarly accurate representations of the intermolecular
interactions. With this new parametrization, simulations using the
TIP4P model can be carried out to determine possible changes for the
structure and spectroscopy when using a water model different from
TIP3P. This, however, is beyond the scope of the present work.
The present work combines state-of-the art electrostatic models with
established empirical energy functions to correctly describe
thermodynamic properties of components used in eutectic mixtures
consisting of KSCN, acetamide and water. The refined models are then
used to characterize structural data (radial distribution functions),
measured spectroscopic properties from 2-dimensional infrared
spectroscopy, and thermodynamic properties (viscosity). It is found
that electrostatic models based on distributed charges (M1 and
M2) perform better than a model using multipoles (M0) to
reproduce the magnitudes of the FFCF decay times for different water
contents as seen in experiments. For the radial distribution functions
it should be noted that they are not direct experimental observables
but rather fitted to reproduce neutron scattering data. Given the
performance of the models, in particular M1 and M2 on the
measured thermodynamic and spectroscopic data it is anticipated that
the computed atom-atom radial distributions functions from simulations
are even more reliable than those determined from experiment.
In summary, the parametrization strategies pursued in the present work
successfully combined electronic structure methods, comparison with
thermodynamic reference data and application to measured structural
and spectroscopic properties. Adjustment of both, electrostatic
contributions and LJ-parameters was found to be important. The present
work suggests that distributed charge models are preferable over
multipole-based representations of the electrostatics for the system
compositions considered here. Application of the procedures outlined
here to other, related mixtures and additional observables will
provide further insight into generalizability and scope for
improvements of this approach.
Acknowledgment
This work has been financially supported by the Swiss National Science
Foundation (NCCR MUST, 200020_219779, 200021_215088 to MM), the
University of Basel (to MM) by European Union’s Horizon 2020 research
and innovation program under the Marie Skłodowska-Curie grant
agreement No 801459 -FP-RESOMUS (to KT).
Supporting Information
The supporting material includes Table S1 to S6 and Figures S1 to S4.
Data Availability
Relevant data for the present study are available at https://github.com/MMunibas/DES2.

### v2

@@
-Force Fields for Deep Eutectic Mixtures: Application to Structure and 2D-Infrared Spectroscopy
+Force Fields for Deep Eutectic Mixtures: Application to Structure, Thermodynamics and 2D-Infrared Spectroscopy
 Abstract
@@
 This is not the case for model M0.
-si University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
+University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
 University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
@@
 theory. On the other hand, empirical energy functions (“force
-fields”) are computationally advantageous to evaluate. Propagating
-the dynamics based on Newton’s equations of motion allows to reach
-long time scales for large molecular systems, e.g. sub-microsecond
-for atoms.2 Such molecular dynamics (MD)
-simulations have been used to investigate processes ranging from
-protein folding,3 ligand binding,4
-crowding in cellular environments5 to characterizing
-spectroscopic properties of solutes and peptides and reactions in the
-gas phase and in solution.6 However one of the
-challenges remains to develop energy functions that retain the
-precision of the QM methods they are often based on.
+fields”) are computationally advantageous to evaluate and are being
+used in molecular dynamics (MD) simulations which solve the
+Newtonian equations of motions. MD simulations have been used to
+investigate processes ranging from protein folding,2
+ligand binding,3 crowding in cellular
+environments4 to characterizing spectroscopic
+properties of solutes and peptides and reactions in the gas phase and
+in solution.5 However one of the challenges remains
+to develop energy functions that retain the precision of the QM
+methods they are often based on.
 One- or multi-dimensional vibrational spectroscopy is a powerful means
 to characterize the structural dynamics of complex
-systems.7, 8 The molecular-level understanding
+systems.6, 7 The molecular-level understanding
 and interpretation of the spectral signatures for such systems is
 greatly aided by a combined simulation/experiment
-approach.9, 10 The atomistic simulations
+approach.8, 9 The atomistic simulations
 often rely on energy functions that parametrize the bonded and
@@
 interactions together with additional, more purpose-tailored
-terms.11 For applications to spectroscopy the
-chemical bonds (stretching vibration) need to be described somewhat
-more realistically to account for mechanical anharmonicity for which
-Morse-oscillators are often sufficient because experiments at ambient
-temperatures are not sensitive to highly vibrationally excited
-states. Nevertheless, there is scope to use more accurate
-representations, for example based on machine learning-type
-approaches, in particular if reference data from high-level electronic
-structure calculations are available. For the nonbonded interactions -
-which include electrostatic and van der Waals terms - more
-physics-based models that go beyond the standard representations have
-been developed.
+terms.10 For applications to vibrational
+spectroscopy, going beyond the harmonic approximation to describe
+chemical bonds is necessary to include effects of mechanical
+anharmonicity. Such improvements can, e.g., be achieved through the
+use of machine learning-based approaches, in particular if
+reference data from high-level electronic structure calculations are
+available.11, 5, 12, 13, 14
+For the nonbonded interactions - which include electrostatic and van
+der Waals terms - more physics-based models that go beyond the
+standard representations have been developed.
 The first-order treatment of the electrostatic interaction is based on
@@
 describing anisotropic contributions to the charge
-density.12 Including higher-order atomic multipoles
+density.15 Including higher-order atomic multipoles
 improves the accuracy but at the expense of increased computational
 cost and implementation
-complexity.13, 14, 15, 16
+complexity.16, 17, 18, 19
 Accounting for polarizability is another contribution that has been
 recently included in empirical force fields and shows much promise for
-further improvements of the computational models.17 From
+further improvements of the computational models.20 From
 an empirical force field perspective the van der Waals interactions
@@
 improved representations are the buffered 14-7
-parametrization18 and modified combination
-rules19, 20
-Deep eutectic mixtures (DES) are electrostatically demanding systems.
+parametrization21 and modified combination
+rules22, 23
+Deep eutectic solvents (DESs) are multicomponent mixtures
+comprising hydrogen bond acceptors (HBAs) and hydrogen bond donors
+(HBDs) at particular molar
+ratios.24, 25, 26 In
+general, DESs are characterized by a pronounced depression of their
+melting point compared to their components and remain in the liquid
+phase over a wider temperature range.27 If the
+mixtures contain ions, the intermolecular interactions involve
+pronounced electrostatic contributions which is also - in part - due
+to crowding. The particular mixture considered here consists of
+water, acetamide and KSCN which is present as solvated K+ and
+SCN- (thiocyanate) ions.28, 29 Acetamide
+forms low-temperature eutectics with a wide range of inorganic salts
+and the resulting non-aqueous solvents have a high ionicity. Such
+mixtures have also been recognized as excellent solvents and molten
+acetamide is known to dissolve inorganic and organic compounds. The
+SCN- anion is an ideal spectroscopic probe because the CN-stretch
+vibration absorbs in an otherwise empty region of the infrared
+region. Previous work has taken advantage of these spectral features to
+probe the
+effect of water addition to urea/choline chloride and in
+acetamide/water mixtures.30, 9
+Traditionally, energy functions for DES have been optimized by
+starting from a conventional energy function and readjusting
+specifically partial charges and in
+some cases van der Waals parameters.31, 32, 33, 34, 35, 36, 37, 38
 The present work aims at parametrizing atomistic force fields using
@@
 Molecular dynamics simulations were performed using the CHARMM
-program.21 The molar composition of the systems
+program.39 The molar composition of the systems
 was changed by modifying the number of water and acetamide molecules
@@
 In total, 5 independent random initial configurations for each of the
-9 system compositions were set up using PACKMOL.24
+9 system compositions were set up using PACKMOL.42
 Following 100 ps of heating and equilibrium simulation, each,
@@
 using the leap-frog integrator and a Hoover
-thermostat.25 For each system composition a total of
-ns were sampled.
+thermostat within the extended system constant pressure and
+temperature algorithm as implemented in CHARMM.43, 44
+The mass of the pressure piston and piston collision frequency were
+set to u and 5 ps-1, respectively, and the mass of the
+thermal piston to kcal/mol ps2.
+For each system composition a total of ns were sampled.
 2.2 Inter- and Intramolecular Interactions
@@
 parametrizations based on the all-atom force field CHARMM36
-(CGenFF)26, the corresponding TIP3P water
-model27 to be used together with CGenFF, and
+(CGenFF)45, the corresponding TIP3P water
+model46 to be used together with CGenFF, and
 optimized parameters for the ions. Two sets of nonbonding
@@
 assigned atom charge of were adopted from previous
-work28, 29 and acetamide was described by
+work47, 48 and acetamide was described by
 CGenFF. Dedicated parametrizations for bonding and nonbonding terms
@@
 anion (SCN-) was a reproducing kernel Hilbert space
-(RKHS)30, 31 representation based on ab
+(RKHS)49, 50 representation based on ab
 initio data at the PNO-LCCSD(T)-F12/aug-cc-pVTZ-F12 level of theory
-using the MOLPRO program package.32, 33 The
+using the MOLPRO program package.51, 52 The
 geometry of SCN- was described in terms of Jacobi coordinates
@@
 through the RKHS representation for SCN- along Jacobi distances
-and for different angles are shown in Figure S1.
+and for different angles are shown in Figure
+S1.
 Electrostatic Interactions: For the electrostatic interactions
@@
 flexible minimally distributed charge model
-(fMDCM).34 For this, 8 point charges were
+(fMDCM).53 For this, 8 point charges were
 distributed around the SCN- atoms with positions within the local
@@
 geometries were determined at the M06-2X/aug-cc-pVTZ level of theory
-using the Gaussian suite of programs.35 The grid
+using the Gaussian suite of programs.54 The grid
 contained 289 geometries, including 32 SCN valence angles (]) and 3 bond lengths for the SC ( Å) and CN ( Å) separations.
@@
 according to the optimization scheme presented in
-Refs. 36, 37. Starting from this model as
+Refs. 55, 56. Starting from this model as
 an initial guess for the polynomial parameters, the quasi-Newton BFGS
-minimization method38 was used to optimize
+minimization method57 was used to optimize
 the polynomial parameters by minimizing the RMSE between reference and
@@
 linear equilibrium structure of SCN-. Following the
-literature,39 the diagonal elements of computed
+literature,58 the diagonal elements of computed
 polarizability for SCN- in the gas phase
@@
 anions from gas phase to water solution by between 5% and 34%,
-respectively.40, 41 Thus, a reduced
+respectively.59, 60 Thus, a reduced
 polarizability of Å3 was equally distributed to the atoms
@@
 In addition to the new fMDCM model, a multipolar
-(MTP)14, 42, 43 representation of
-SCN- used in previous work was used as well.10
+(MTP)17, 61, 62 representation of
+SCN- used in previous work was used as well.9
 The electrostatic atomic moments up to the quadrupole moment were
@@
 also requires to readjust the Lennard-Jones
-parameters.44 Three models were considered in the
+parameters.63 Three models were considered in the
 following for SCN-. Model (M0) was based on MTP with LJ
-parameters originally from Bian et al.28 and
+parameters originally from Bian et al.47 and
 scaled to best reproduce measured thermodynamic properties. Model M1 used fMDCM electrostatic instead of MTP and initial LJ
-parameters from Tesei et al.,29 and model M2
+parameters from Tesei et al.,48 and model M2
 was a combination of fMDCM and optimized LJ parameters with respect to
@@
 cation the LJ parameters were those from the respective parameter set
-in the literature (Ref. 28 for M0 and M2
-and Ref. 29 for M1) and for model M2 the
-LJ parameters from Bian and coworker were chosen for
-K+.28
+in the literature (Ref. 47 for M0 and M2
+and Ref. 48 for M1).
+For model M2, also based on the fMDCM representation for
+the SCN- anion, a different approach was taken. Here, the LJ
+parameters and were adjusted to best
+match counterpoise-corrected64 interaction energies
+from electronic structure calculations. The systems considered
+included one SCN- anion surrounded by (i) 16 water molecules;
+(ii) 14 water molecules with one additional SCN-, (iii)
+14 water molecules with one K+ ion; (iv) 12 water molecules
+with both one additional SCN- and K+ ion. For each type
+of system (i) to (iv) 50 independent conformations were extracted
+randomly from previous MD simulations.9
+Reference interaction energies were then determined at the
+M06-2X/aug-cc-pVTZ level of theory using Gaussian
+program.54 Higher levels of quantum chemical theory,
+such as coupled cluster energies, are not feasible due to the
+unfavourable scaling of the computations with system and basis set
+size.
 2.3 Analysis
-Hydration free energies for the SCN- anion in water solvent were computed from thermodynamic integration.9 One SCN- anion was sampled in the gas phase and in pure water.
-The condensed-phase simulations were carried out in the NpT ensemble with 997 water molecules (cubic box size Å3).45, 9 Simulations with 24 different coupling parameters for electrostatic and vdW interactions, respectively, were performed for each a gas phase and solvated SCN-.
+Hydration free energies for the SCN- anion in water solvent were computed from thermodynamic integration.8 One SCN- anion was sampled in the gas phase and in pure water.
+The condensed-phase simulations were carried out in the NpT ensemble with 997 water molecules (cubic box size Å3).65, 8 Simulations with 24 different coupling parameters for electrostatic and vdW interactions, respectively, were performed for each a gas phase and solvated SCN-.
 Initial conditions for these simulation were taken from an unbiased simulation, equilibrated for 50 ps with the respective coupling parameter and run for another 150 ps for statistical sampling.
@@
 due to the 1-2 and 1-3
-nonbonded interaction exclusion.46 Therefore, only
+nonbonded interaction exclusion.66 Therefore, only
 and
@@
 500 ps.
-For computing the frequency fluctuation
-correlation function (FFCF), the frequency trajectories
-for each oscillator are required. These were determined from an
-instantaneous normal mode (INM)47 analysis for the CN
-vibrational frequencies for all SCN- ions
-on snapshots of the trajectories every fs along
-the first 2 ns of each production simulation (10 ns in total for
-each system composition).
-For the INM analysis the structure of every SCN- ion was optimized
-while freezing the position of all remaining atoms in the system,
-followed by a normal mode analysis using the same force field that was
-employed for the MD simulations. Previously, such an approach has been
-validated for N in solution by comparing with rigorous quantum
-bound state calculations.48
+For computing the frequency fluctuation correlation function (FFCF),
+the frequency trajectories for each oscillator are
+required. These were determined from an instantaneous normal mode
+(INM)67 analysis for the CN vibrational frequencies
+for all SCN- ions on snapshots of the trajectories every
+fs along the first 2 ns of each production simulation (10 ns
+in total for each system composition). For the INM analysis the
+structure of every SCN- ion was optimized while freezing the
+position of all remaining atoms in the system, followed by a normal
+mode analysis using the same force field that was employed for the MD
+simulations. Previously, such an approach has been validated for
+N in solution by comparing with rigorous quantum bound state
+calculations.11
 From the frequency trajectory for each oscillator the FFCF was determined which contains information on relaxation time scales corresponding to the solvent dynamics around the solute.
-The FFCFs were fit to an empirical expression
-using an automated curve fitting tool from the SciPy
-library.49 Here, , , and
+The FFCFs were fit to an empirical expression68, 8
+using an
+automated curve fitting function (scipy.optimize.curve_fit)
+from the SciPy library using the default trust region reflective
+algorithm.69
+Here, , , and
 are the amplitudes, decay time scales, phase and
@@
 2.4 2D-IR Spectroscopy
-The experimental data shown in this paper (Figure 7) originate from the same raw data as those of Ref.
-10, this time extracting the FFCF from the 2D-IR spectra.
-In brief, a 2D IR spectrometer in pump-probe with scatter suppression by quasi-phase-cycling has been used to obtain the raw data.50 A 50%/50% mixture of two isotopologues of the salt, KS12C14N and KS13C15 has been investigated.
-Solutions were prepared with constant total salt concentration of 4.4 M, and the molar ratio of D2O vs acetamide varied from 0% to 100% in steps of 10%.
-To extract the essential information from the 2D IR spectra, i.e., the FFCF, each 2D IR peak was modelled as a correlated 2D-Gaussian function:
-where is an amplitude, the center frequencies along
-frequency axis , the frequency width, and a
-correlation coefficient.51 In total 8 such peaks were
-defined for a 2D IR spectrum, with many parameters set equal to
-minimize their total number in the global fitting. That is, two pairs
+The experimental data shown in this paper (Figure 7)
+originate from the same raw data as those of
+Ref. 9, this time extracting the FFCF from the
+2D-IR spectra. Technical details of the 2D IR spectrometer used
+to obtain the raw data are described in Ref. 70. A
+50%/50% mixture of two isotopologues of the salt, KS12C14N
+and KS13C15N has been investigated. Solutions were
+prepared with constant total salt concentration of 4.4 M, and the
+molar ratio of D2O vs acetamide varied from 0% to 100% in steps
+of 10%.
+Very generally speaking, the 2D IR spectrum of two coupled modes consist of two diagonal contributions for each mode, and two cross peak contributions that appear when energy transfer occurs between both modes.
+Furthermore, each one of these contributions consist of a pair of a negative and positive peak, i.e., ground state bleach and excited state absorption, respectively.
+Hence, eight peaks exist in total.7 To extract the essential information from the 2D IR spectra, i.e., the FFCF, each one of these 2D IR peaks was modelled as a correlated 2D-Gaussian function:
+where is an amplitude, the center frequencies along frequency axis , the frequency width, and a correlation coefficient.71 In the limiting case of , the 2D Gaussian is uncorrelated and a circular peak would appear.
+For the 2D Gaussian is fully correlated with the peak stretched along the diagonal and widths in the diagonal direction and 0 in the anti-diagonal direction.
+To minimize the total number of parameters in the global fitting, many of the parameters for modeling the eight peaks as Gaussian functions were constrained to be equal.
+That is, two pairs
 of diagonal peaks were defined for the two isotopologues with
@@
 two parameters describe the identical interaction of the ions with the
-solvent environment. Each diagonal peak consists of ground state
-bleach (GB)/stimulated emission (SE) as well as excited state
-absorption (EA), where the latter has been down-shifted in the
+solvent environment. For each diagonal contribution, the ground state
+bleach was positioned on the diagonal (i.e., ), while the excited state absorption has been down-shifted in the
 -direction by the anharmonicity
@@
 the two contributions to each diagonal peak were assumed to be the
-same. In addition, two cross peak pairs were introduced, each again
-consisting of GB/SE and EA contributions, for which center frequencies
-and width were assumed to be the same as for the corresponding
-diagonal peaks along the two frequency axes. The amplitudes of the
-two cross peak pairs were set to be the same, and was assumed
-for the correlation, since cross-relaxation randomizes the
-frequency.
+same. The center frequencies and widths of the cross peak pairs were assumed to be the same as for the corresponding diagonal peaks. The
+amplitudes of the two cross peak pairs were set to be the same,
+and was assumed for the correlation, since cross-relaxation
+randomizes the frequency.
 The property of interest here is the frequency correlation of the diagonal peaks as a function of the waiting time between the pump-pulse pair and the probe pulse.
-For times larger than a few 100 fs, reflects the FFCF and can be compared to results from MD simulations after proper scaling, while dynamics faster than this is hidden in the homogeneous dephasing time .8
+It evolves from ideally 1 at to 0 for , however, dynamics faster than few 100 fs is hidden in the homogeneous dephasing time and cannot be observed.7 Beyond a few 100 fs, reflects the normalized FFCF and can be compared to results from MD simulations after proper scaling.
 3 Results
@@
 ( g/cm3) at the equivalent KSCN molality
-of mol/kg52 and the
+of mol/kg72 and the
 hydration free energy . For SCN- an
 estimated value of kcal/mol is
-available53 which compares with related anions such as
+available73 which compares with related anions such as
 HS- ( kcal/mol), N ( kcal/mol)
-54, or CN- ( kcal/mol)55.
+74, or CN- ( kcal/mol)75.
 For model M0 using the MTP electrostatic model, the
-literature28 LJ parameters of the
+literature47 LJ parameters of the
 SCN- atoms were scaled by multiplying with a factor . Using these scaled parameters, the
@@
 Model M1 (fMDCM electrostatics) was based on LJ parameters that
-had been used for modeling aqueous SCN- solutions.29
+had been used for modeling aqueous SCN- solutions.48
 To adjust the literature LJ parameters to the fMDCM electrostatic
@@
 computed density g/cm3 compares well
-with the experimentally measured52 density
+with the experimentally measured72 density
 g/cm3 and the computed hydration free
@@
 range of the experiments.
-For model M2, also based on the fMDCM representation for the
-SCN- anion, a different approach was taken. Here, the LJ parameters
-and were adjusted to best match
-counterpoise-corrected interaction energies from electronic structure
-calculations. The systems considered included on SCN- anion
-surrounded by (i) 16 water molecules; (ii) 14 water molecules with
-either 1 SCN- or (iii) 1 K+ ion; (iv) 12 water molecules with
-both 1 SCN- and K+ ion. For each type of system (i) to (iv) 50
-independent conformations were extracted randomly from previous MD
-simulations.10 Reference interaction energies were
-then determined at the M06-2X/aug-cc-pVTZ level of theory using
-Gaussian program.35 Higher levels of quantum chemical
-theory, such as coupled cluster energies, are not feasible due to the
-unfavourable scaling of the computations with system and basis set
-size.
 The interaction energy consists of electrostatic contributions using
@@
 Figure 2 (red lines). Closest agreement with
-experiment is found for which yields
+experiment was found for , very close to
+(no modifications in the van der Waals parameters), which yields
 g/cm3 and kcal/mol. Hence, fitting the
@@
 relaxation and hydration free energies of CN- and
-OH-.9
+OH-.8
 3.2 Structure and Ordering of the Mixture
@@
 neutron diffraction measurements of aqueous KSCN
-solution.56 It is important to stress that such pair
+solution.76 It is important to stress that such pair
 distribution functions are not measured directly from experiments but
@@
 amplitude which is written as a weighted average over pair
-distribution functions.56 Figure 4
+distribution functions.76 Figure 4
 reports SCN-–water (top) and SCN-–SCN- (bottom) pair
@@
 between simulations (dot-dashed black line) and EPSR results (solid
-black line) are possible.56 For model M0, the
+black line) are possible.76 For model M0, the
 overall shape of agrees reasonably well with the
@@
 maximum is overestimated for M1 but quite well described by M2. With decreasing water content the height of the first peak
-increases - i.e. the ion recruits water molecules as the water density
-decreases. The shift of the first maximum for model M2 relative
-to model M1 is consistent with the larger values
-for the S- and N-atoms for M2 compared with M1, see Tables
-S3 and S4.
+increases. In other words, with decreasing water density the ion
+recruits water molecules. The shift of the first maximum for model
+M2 relative to model M1 is consistent with the larger
+values for the S- and N-atoms for M2 compared with
+M1, see Tables S3 and
+S4.
 For the C–OW and S–OW pair correlation functions
@@
 either shifted to longer separations or a much too pronounced
-maximum. Compared with that, models M1 and M2 correctly
-capture the peak height, which is a proxy for the interaction
-strength, whereas the position of the first maximum is in general
-shifted to shorter separations. This, however, changes with decreasing
-water content for which the S–N and N–N pair correlation functions
-shift to longer separations. It should be noted that the peak position
-of the anion-anion pair distribution functions in simulations can be
-steered by adjusting the LJ parameter of the SCN-
-atoms individually. However, the procedure followed here was not to
-match pair distribution functions as the LJ parameters of
-SCN- are at first either taken from literature (M0, M1)
-or fit to ab initio data (M2) and later scaled
-collectively to best match the density of an aqueous KSCN system or
-hydration free energy of SCN- in water.
-The comparison of the between atoms of different SCN- anions
-in simulations using the three models for the intermolecular
-interactions are shown in Figures 4D to F. Peak
-intensities for SCN- atom pair in different water/acetamide solvent
-mixture remains either similar (M1) or increase for larger
-acetamide content (M2). For simulations with multipoles (model
-M0), the peak positions are generally closer those reported from
-experiments but with higher intensities which are highest for pure
-water solvent and decreases for larger acetamide content in the
-water/acetamide solvent mixture.
-Another view on the structural SCN- radial-angular pair
-distribution is shown in Figure 5 along the C-C bond
-distances and (CN)-(CN) bond angles (similar to SCN- dipole
-vectors). Even after adjusting the LJ parameters for M0 to yield
-improved thermodynamics, the MTP-based model lead to mainly angled
-SCN- pairs in pure acetamide and solvent mixtures with large
-acetamide content. This changes to predominantly parallel and, partly,
+maximum. Compared with that, models M1 and M2
+correctly capture the peak height for and , which is a proxy for the interaction strength. The
+position of the first maximum, however, is at larger separation as
+seen for model M0. Opposite to the reference EPSR results,
+the position of the first maximum in is at larger
+separation than in due to a higher concentration of
+negative charge on the N- rather than the S-atom in SCN-
+according to the fitted fMDCM ( e
+vs. e) and MTPL ( e vs. e) models.
+It should
+be noted that the peak position of the anion-anion pair distribution
+functions in simulations can be steered by adjusting the LJ parameter
+of the SCN- atoms individually. However, the
+procedure followed here was not to match pair distribution functions
+as the LJ parameters of SCN- are at first either taken from
+literature (M0, M1) or fit to ab initio data (M2) and later scaled collectively to best match the density of an
+aqueous KSCN system or hydration free energy of SCN- in water.
+Peak intensities for SCN- atom pairs in different water/acetamide
+solvent mixture remains either similar (M1) or increase for
+larger acetamide content (M2). For simulations with multipoles
+(model M0), the peak positions are generally closer those
+reported from experiments but with higher intensities which are
+highest for pure water solvent and decreases for larger acetamide
+content in the water/acetamide solvent mixture.
+Figure 5 reports radial-angular pair distribution
+functions for SCN-/SCN- pairs along the C-C bond distances
+and (CN)-(CN) bond angles (related to SCN- dipole vectors). Even
+after adjusting the LJ parameters for M0 to yield improved
+thermodynamics, the MTP-based model lead to mainly angled SCN-
+pairs in pure acetamide and solvent mixtures with large acetamide
+content. This changes to predominantly parallel and, partly,
 anti-parallel SCN- pair alignments with increasing water content
 all of which is consistent with results from simulations using the
-unmodified MTP-based model.10 In simulations using
-fMDCM for the electrostatics (models M1 and M2), the
+unmodified MTP-based model.9 In simulations
+using fMDCM for the electrostatics (models M1 and M2), the
 angular pair SCN- distribution functions depend less strongly on
-solvent mixtures and are close to squared cosine angle distributions
-which are statistically expected for weakly or non-interacting
-particles. This can be even better seen in Figure 6
-which reports the 2-dimensional projections of the distributions shown
-in Figure 5 using model M2.
+solvent mixtures and are close to a sine-angle distribution
+which is the probability to find a particle at a certain
+angle on a sphere ().
+This can be even better seen in
+Figure 6 which reports the 2-dimensional projections
+of the distributions shown in Figure 5 using model M2.
 Contrary to M0, the probability distributions do not change so
@@
 higher-probability regions at large C–C separations Å
-that appear for acetamide as the solvent disappear as the water
-content increases.
+that are present with acetamide as the solvent disappear with
+increasing water content.
 3.3 Dynamics of the Mixtures
@@
 bi-exponential fit is a meaningful model applicable to all
-mixtures. Assuming a single exponential significantly deteriorates the
-fit. To allow direct comparison between experiment and simulations,
-the experimental data was fit to a double-exponential decay because
-for the computed FFCFs a second, short-time decay (to which the
-experiments are insensitive) needs to be included for fitting the full
-time range of the data, see Figures S2A to C. The
-set of optimized parameters of the bi-exponential fit function are
-shown in Figure S3 and Table S6.
-Qualitatively, models M1 and M2 are rather successful in
-capturing the experimentally observed FFCFs and their dependence on
-the amount of water in the solvent mixture whereas model M0 is
-not. Overall, the decay from simulations using M0 is too
-slow. However, the relative amount by which the time constant changes
-between 0% and 100 % water is considerably closer to the
-experimentally observed change than in the other two models.
+mixtures. Assuming a single exponential significantly
+deteriorates the fit. To allow direct comparison between experiment
+and simulations, the experimental data was also fit to a
+double-exponential. The set of optimized parameters of the
+bi-exponential fit function are shown in Figure S3
+and Table S6. Qualitatively, models M1
+and M2 capture the experimentally observed FFCFs and their
+dependence on the amount of water in the solvent mixture whereas
+model M0 does not. Overall, the decay on the longer
+time scale from simulations using M0 is too slow. However,
+the relative amount by which the time constant changes between 0% and
+100 % water is considerably closer to the experimentally observed
+change than in the other two models.
 A more quantitative analysis is based on the decay times of the FFCFs
@@
 using different protocols and
-approaches57, 58, 59, 60, 61, 62, 63, 64
+approaches31, 32, 33, 34, 35, 36, 37, 38
 Most common to these efforts is to start from a conventional energy
@@
 77 adapted for particular
-applications66 and to further readjust
-parameters.57, 63 The adjustment is usually
-to scale partial charges62 and in some cases
+applications78 and to further readjust
+parameters.31, 37 The adjustment is usually
+to scale partial charges36 and in some cases
 van-der-Waals parameters to reproduce observed properties such as
-diffusivities, viscosities or the densities.63 In terms
+diffusivities, viscosities or the densities.37 In terms
 of using diffusivities for the parametrization from explicit MD
@@
 m2/2 vs.
-m2/s).61, 57, 67 Thus the
-determination of the values typically require a set of ns simulations (compared with 5 ns for water67)
+m2/s).35, 31, 79 Thus the
+determination of self-diffusion coefficients typically requires a set of ns simulations (compared with 5 ns for water79)
 which would make an explicit parametrization extremely expensive.
 Alternatively, there have been efforts to develop models based on
-symmetry adapted perturbation theory (SAPT)61 which
+symmetry adapted perturbation theory (SAPT)35 which
 were refined by comparing with first principles MD
@@
 Padua (CL&P) force field and its polarizable variant
-CL&Pol.58 Recent fine-tuning efforts used structural
-data.59 The present status of computational approaches
+CL&Pol.32 Recent fine-tuning efforts used structural
+data.33 The present status of computational approaches
 for deep eutectic solvents has been recently
-reviewed.64
+reviewed.38
 In the present work two alternative procedures were followed, both
@@
 followed to best describe again experimental liquid densities and
-hydration free energies as for M0 and M1.
+hydration free energies as for M0 and M1. This
+strategy follows common practices in empirical energy function
+development in that the electrostatic parameters usually originate
+from high-level quantum chemical calculations whereas the
+van-der-Waals parameters are adjusted with respect to, for example, interaction
+energies and hydration free energies.80
 Interestingly, the optimal scaling parameter for model M2 to
@@
 )) needed to be increased by % for hydration free energies consistent with
-experiment.9
+experiment.8
 Decomposition of a molecular ESP into atomic distributed multipoles
@@
 transfer between neighboring SCN-, which was measured
-previously,10 is affected by intermolecular
+previously,9 is affected by intermolecular
 coupling and structural alignment. The more pronounced anisotropy of a
@@
 compared with the earlier employed
-parametrization10 do not differ appreciably
-(compare with Figure 5 in Ref.10)
+parametrization9 do not differ appreciably
+(compare with Figure 5 in Ref.9)
 Using the three interaction models the viscosities for varying
@@
 are the upper triangular elements of the modified stress tensor
-.68, 63 Due to the
+.81, 37 Due to the
 strong intermolecular interactions and high viscosities, converging
@@
 two. Experiments for the water-free system report
-mPas at 303 K.69 Simulations with M1 and
+mPas at 303 K.82 Simulations with M1 and
 M2 are consistent with such values. More importantly, however,
@@
 various limitations when comparing experimentally measured quantities
-with those from simulations.70 As was already
+with those from simulations.83 As was already
 mentioned, the reason to use the TIP3P model was consistency with the
@@
 model, MD simulations were carried out using model M2 but with
-the TIP4P water model.27 Again, hydrated
+the TIP4P water model.46 Again, hydrated
 SCN- clusters were extracted, reference interaction energies
@@
 for TIP3P, see Figure 3. The results are reported in
-Figure 9. It is found that the two water models yield
-similarly accurate representations of the intermolecular
-interactions. With this new parametrization, simulations using the
-TIP4P model can be carried out to determine possible changes for the
-structure and spectroscopy when using a water model different from
-TIP3P. This, however, is beyond the scope of the present work.
-The present work combines state-of-the art electrostatic models with
-established empirical energy functions to correctly describe
-thermodynamic properties of components used in eutectic mixtures
-consisting of KSCN, acetamide and water. The refined models are then
-used to characterize structural data (radial distribution functions),
-measured spectroscopic properties from 2-dimensional infrared
-spectroscopy, and thermodynamic properties (viscosity). It is found
-that electrostatic models based on distributed charges (M1 and
-M2) perform better than a model using multipoles (M0) to
-reproduce the magnitudes of the FFCF decay times for different water
-contents as seen in experiments. For the radial distribution functions
-it should be noted that they are not direct experimental observables
-but rather fitted to reproduce neutron scattering data. Given the
-performance of the models, in particular M1 and M2 on the
-measured thermodynamic and spectroscopic data it is anticipated that
-the computed atom-atom radial distributions functions from simulations
-are even more reliable than those determined from experiment.
-In summary, the parametrization strategies pursued in the present work
-successfully combined electronic structure methods, comparison with
-thermodynamic reference data and application to measured structural
-and spectroscopic properties. Adjustment of both, electrostatic
-contributions and LJ-parameters was found to be important. The present
-work suggests that distributed charge models are preferable over
-multipole-based representations of the electrostatics for the system
-compositions considered here. Application of the procedures outlined
-here to other, related mixtures and additional observables will
-provide further insight into generalizability and scope for
-improvements of this approach.
+Figure 9 and Table S4. It is
+found that the two water models yield similarly accurate
+representations of the intermolecular interactions. With this new
+parametrization, simulations using the TIP4P model can be carried out
+to determine possible changes for the structure and spectroscopy when
+using a water model different from TIP3P. This, however, is beyond the
+scope of the present work.
+Another future extension of the present work is to use kernel-based MDCM (kMDCM) to capture conformationally dependent changes in the ESP in particular for molecules larger than triatomics.84
+In summary, the parametrization strategies pursued in the present
+work successfully combined electronic structure methods, comparison
+with thermodynamic reference data and application to measured
+structural and spectroscopic properties. Adjustment of both,
+electrostatic contributions and LJ-parameters was found to be
+important. The present work suggests that distributed charge models
+are preferable over multipole-based representations of the
+electrostatics for the system compositions considered
+here. Application of the procedures outlined here to other, related
+mixtures and additional observables will provide further insight
+into generalizability and scope for improvements of this
+approach.
 Acknowledgment